Р с D B 40 KN 40 KN 0.6 m 1.8 m 0,6 m

For the beam shown, calculate the reactions at A and B, and then draw the shear and Bending-Moment diagrams. Consider the loading P= 31 kN/m.
Answer #1

Answer: P=31 kN/m 0.6m 1.8m 0.6m. R | Ro YORN yoki 41 (27) = 0 RA+ R$ = 31x1.8+ 40+40 RA+Ro = 135.8KN . EMA=0 R2x3 () + 40x0.(ii) Section 00:- P=31 kNlm - Vx - 3/4-40+ 67.920 Voc 0.6m bc = 27.9-313 Ve os Just right of c? = 27.9 KN Mx YOKN Re: 7 9k Vx67.9KN 27.9kW A Vx=0 0.9m 14.-27.9KN -6.9KN Figure: Shear force diagram (SFD) Mmax = 53.7KN-m 40.74KN-m 40.74 KNM B re Figure